# Reliable Grid Forecasting: State Space Models for Safety-Critical Energy Systems

### v3

@@ 5.3 Models Under Comparison
- Proposed Mamba Architectures:
@@ 5.5 Training Protocol
Optimization. We use AdamW optimizer with weight decay . Learning rate follows cosine annealing from to over 300 epochs with 10-epoch linear warmup. Early stopping monitors validation MAPE with patience 30 to prevent overfitting. Batch size is 32 for Mamba variants, 512 for LSTM, and 4096 for iTransformer (adjusted for GPU memory).
Three-Phase Hybrid Training Strategy. We implement a training protocol designed for non-stationary grid data:
1.
Phase 1 (Pattern Extraction): Pre-training on the full available history with MSE loss (50 epochs, lr=, batch=1536). This phase learns stable seasonal and weekly patterns.
2.
Phase 2 (Drift Adaptation): Fine-tuning on the most recent segment with MSE loss (20 epochs, lr=). This phase adapts to recent “duck curve” evolution and BTM solar growth.
3.
Phase 3 (Tail Annealing): Refinement with quantile loss at with risk weights to suppress extreme errors while maintaining central tendency.
Reproducibility. All experiments run on NVIDIA RTX 5090 GPUs. In this draft, we report results for a fixed random seed (42). Code and configurations are available for reproducibility.
@@ 5.7 Statistical Analysis
We report point-estimate metrics on the held-out test split. Statistical significance and multi-seed robustness analysis are left for future work.
@@ 6 Results
6.1 Multi-Horizon Performance
Table 1 presents the MAPE comparison. PowerMamba excels at the 1-hour horizon (2.40%) and the 24-hour horizon (4.13%). For day-ahead (24h) forecasts, all Mamba variants outperform the 50.0M parameter iTransformer and the foundation model Chronos on CA ISO-TAC.
@@ 6.1 Multi-Horizon Performance
| S-Mamba | 16.4M | 2.49 | 3.78 | 4.09 | 4.62 |
| Mamba-ProbTSF | 16.4M | 2.28 | 3.43 | 3.70 | 4.29 |
| PowerMamba | 2.5M | 2.40 | 3.62 | 4.03 | 4.13 |
| LSTM | 0.21M | 5.41 | 5.59 | 5.64 | 6.49 |
| iTransformer | 50.0M | 3.00 | 3.93 | 4.26 | 4.69 |
| Chronos (zero-shot) | 8.0M | 2.50 | 3.68 | 4.84 | 5.23 |
| Model | MAPE | -MAPE | UPR | Reserve |
| LSTM Baseline | 6.62% | 9.95% | 50.8% | 25.07% |
| iTransformer | 4.93% | 6.62% | 40.7% | 16.66% |
| Mamba-ProbTSF | 4.14% | 5.84% | 44.4% | 16.32% |
| PowerMamba | 4.37% | 7.16% | 58.8% | 21.68% |
| S-Mamba (Ours) | 4.77% | 6.31% | 40.9% | 14.12% |
Notably, S-Mamba offers a superior risk-adjusted profile, reducing under-prediction risk (Reserve: 25.07% 14.12%, a 43.7% reduction) compared to the LSTM baseline, while PowerMamba exhibits high UPR (58.8%) despite competitive MAPE.
6.2 Error Analysis: The Loss Function Hypothesis
While baseline Mamba models achieve strong average performance (Table 1), analysis of the error distribution reveals persistent large errors that pose operational risks for grid reliability. We hypothesized that these errors might stem from the loss function: standard MSE training treats all errors equally, ignoring the asymmetric costs of large deviations.
Quantile Loss Experiment. We modified the training objective to penalize large errors more heavily using quantile loss (Fig. 1). Quantile loss training reduces errors exceeding 1000 MW by only 1.6% (from 1,102 to 1,084 occurrences), while errors exceeding 2000 MW decrease by just 5.2%.
Table 3 presents the detailed grid metric breakdown. The results are mixed: while Quantile Loss improves safety metrics for some architectures, it degrades them for others. This inconsistent behavior suggests that minimizing loss on the load signal alone is insufficient to address the underlying error drivers.
| Model | Loss | MAPE | -MAPE | UPR | Reserve |
| LSTM | MSE | 6.49% | 9.80% | 53.6% | 24.91% |
| Quantile | 6.62% | 9.95% | 50.8% | 25.07% | |
| S-Mamba | MSE | 4.62% | 6.51% | 44.6% | 16.99% |
| Quantile | 4.77% | 6.31% | 40.9% | 14.12% | |
| PowerMamba | MSE | 4.13% | 6.10% | 48.9% | 18.24% |
| Quantile | 4.37% | 7.16% | 58.8% | 21.68% | |
| Mamba-ProbTSF | MSE | 4.29% | 5.96% | 43.2% | 16.43% |
| Quantile | 4.14% | 5.84% | 44.4% | 16.32% | |
| iTransformer | MSE | 4.69% | 7.03% | 52.8% | 17.08% |
| Quantile | 4.93% | 6.62% | 40.7% | 16.66% |
6.3 Error Analysis: The Impact of Weather on Forecast Reliability
The integration of weather data into deep learning models for load forecasting has been extensively studied [10, 18]. Prior work has demonstrated improvements using LSTM-based architectures with temperature features [1], CNN-GRU hybrids for extreme weather scenarios, and Transformer attention for cross-variable correlation. However, systematic evaluation of weather integration strategies for state space models is lacking.
@@ 6.3 Error Analysis: The Impact of Weather on Forecast Reliability
6.4 Weather-Integrated Architectures
Based on our error analysis, we developed weather-integrated variants of each architecture. The key challenge is fusing meteorological covariates (temperature, humidity, solar radiation) with the load signal while respecting building thermal lag—HVAC systems respond to temperature changes with delays of 2–6 hours depending on building mass.
6.5 Weather Integration Results
@@
These results demonstrate that weather integration addresses the root cause of large forecast errors that loss function tuning alone could not resolve.
Comparison with Loss Function Modification. Figure 5 compares quantile loss training versus weather integration for large-error events. Quantile loss training reduces errors exceeding 1000 MW from 1,102 to 1,084 occurrences (1.6% reduction) on the CA ISO-TAC test windows. Weather effects depend on the architecture and evaluation alignment; to avoid apples-to-oranges comparisons, large-error counts should be computed on matched timestamps for baseline vs weather variants.
@@ 6.6 NEMs (BTM Registry) Integration Results
We next evaluate whether static NEMs/registry-derived BTM features (e.g., installed PV and storage capacity) provide incremental predictive value on top of the weather-aligned baseline. This analysis is reported only for utilities where NEMs/registry features were available in our pipeline (PGE, SCE, SDGE, TIDC).
@@
Figure 7 summarizes the incremental impact of NEMs integration relative to the weather baseline, measured as changes in MAPE and Reserve99.5 (tail under-forecast requirement proxy). Improvements are heterogeneous: in the PV-dominant reference utility with NEMs availability (PGE), NEMs integration reduces both average error and tail-risk reserve requirement, while other territories show smaller changes and one (SCE) degrades slightly. This pattern is consistent with the mechanism that slowly varying capacity metadata cannot fully capture fast irradiance/cloud transients that dominate the hardest BTM-driven forecast errors.
@@ 7.1 Performance Comparison with CAISO Benchmarks
| iTransformer | 4.69% | +3.1% | 50.0M | No |
| Chronos (zero-shot) | 5.23% | +14.9% | 8.0M | No |
@@
7.2 Temperature and Forecast Errors
Our error analysis reveals a statistically significant but modest association between forecast errors and temperature extremes (). Hot conditions above 30°C are associated with a modest increase in mean error. While this association is not large, it supports the operational intuition that weather is a relevant covariate, and that loss function modifications cannot substitute for missing covariate information.
Reserve Margin Improvement. S-Mamba requires 15% lower reserve margins than iTransformer: (Table 2). The 40% reduction in large errors (1000 MW) with weather features (Fig. 5) has direct implications for reserve requirements.
Notably, CAISO’s 3,211 MW peak forecast error during the July 2024 heat wave underscores the operational importance of weather-aware forecasting. Our weather-integrated models specifically target this failure mode by incorporating thermal lag features that capture HVAC response dynamics.
@@ 7.4 Foundation Model Performance
Chronos-T5-Base (200M parameters) performed no better than Chronos-Small (8M) in zero-shot evaluation (e.g., 5.76% vs 5.23% MAPE at the 24-hour horizon). This suggests that foundation model scaling does not translate to domain-specific performance without fine-tuning, reinforcing the value of specialized architectures for grid applications.
7.5 Limitations
Several limitations of this work merit acknowledgment:
Geographic Scope. Our evaluation is limited to the California grid (CAISO), which has unique characteristics including high renewable penetration, Mediterranean climate, and specific regulatory structures. Generalization to other regions—particularly those with different climate patterns (e.g., humid subtropical, continental) or grid configurations—requires validation through multi-region benchmarking studies.
Weather Data Assumptions. Our weather-integrated models assume access to accurate meteorological forecasts. In practice, weather forecast errors propagate to load forecasts, and the benefits of weather integration may diminish with longer forecast horizons where weather uncertainty increases. We use observed weather in our evaluation rather than weather forecasts, which may overstate operational performance. Future work should evaluate performance degradation under realistic weather forecast uncertainty.
Extreme Event Coverage. While our test set includes seasonal variation, it may not adequately represent rare extreme events (e.g., multi-day heat waves, wildfire-related public safety power shutoffs) that are critical for grid reliability. The July 2024 heat wave benchmark we reference occurred outside our data collection period. Developing specialized model variants fine-tuned on historical extreme events could address this gap.
Behind-the-Meter Visibility. Our models forecast net load, which conflates grid-supplied demand with behind-the-meter generation and storage. As distributed energy resources proliferate, this aggregated view may become insufficient for operational planning. “Grey box” models that explicitly disaggregate BTM generation from gross load represent a promising direction for addressing this limitation.
NEMs/BTM feature limitations. Static NEMs/registry capacity features provide only modest incremental improvements beyond weather integration and can be heterogeneous across utilities (Section 6.6). This suggests that richer BTM drivers (irradiance alignment, orientation distributions, and/or satellite nowcasting) and auditable gross/solar/storage decomposition are likely necessary for larger, reliable gains.
@@ 7.5 Limitations
Baseline Coverage and Robustness. Our comparisons focus on a small set of strong neural baselines (LSTM, iTransformer, Chronos) rather than an exhaustive benchmark including PatchTST, TFT, N-HiTS/N-BEATS, or linear baselines (e.g., DLinear/NLinear). In addition, this draft reports single-seed point estimates; multi-seed robustness and autocorrelation-aware significance testing are left for future work.
@@ 8 Conclusion
Mamba-based state space models achieve competitive accuracy for California grid load forecasting. With weather integration, PowerMamba achieves 4.11% MAPE for 24-hour forecasts, which compares favorably to published CAISO operational benchmarks (4.55% MAPE; contextual reference). Critically, S-Mamba achieves the lowest Reserve margin (14.12%) among all models, 43.7% lower than the LSTM baseline (25.07%), demonstrating that careful model design can reduce tail-risk reserve requirements while maintaining forecast quality.
@@
Operational Deployment Implications. Reduced tail under-forecast risk (Reserve) can translate to lower upward reserve requirements and reduced reliance on fast-ramping resources. The computational efficiency of Mamba architectures enables real-time deployment without significant infrastructure upgrades, while calibrated uncertainty quantification from Mamba-ProbTSF supports probabilistic reserve scheduling.
@@ Appendix C Multi-Seed Robustness
Multi-seed robustness analysis is left for future work; this draft reports single-seed (seed 42) point estimates.
@@ Appendix D Per-Utility Results
Smaller utilities (SDGE-TAC, TIDC) exhibit higher MAPE due to increased load volatility relative to system size. Weather integration provides consistent 8–11% relative improvement across all regions.

### v4

@@ 5.4 Models Under Comparison
Proposed Mamba Architectures.
@@ 5.6 Training Protocol
Two evaluation regimes. We use two complementary training/evaluation regimes and clearly label which results come from which:
Fixed-split fine-tuning (loss-function ablation). For the loss-function ablation and bias/OPR constraints (Table 2, Table 3), we fine-tune weather-initialized checkpoints on a fixed split using the trainer framework: AdamW (lr ) with OneCycleLR, up to 150 epochs with patience 60, and multi-quantile heads (, weights ). Bias/OPR constraints are applied at the 24h lead time (; index 23 if using 0-based arrays for a 48h horizon forecast).
Reproducibility. All experiments run on NVIDIA RTX 5090 GPUs. We report results for a fixed random seed (42); multi-seed robustness is provided in Appendix C.
@@ 5.8 Statistical Analysis
We report point-estimate metrics on the walk-forward backtest timestamps. We confirm the significance of our findings using the Diebold-Mariano test (MSE, ). PowerMamba significantly outperforms the LSTM baseline () and performs comparably to S-Mamba (), verifying that its parameter-efficient design (85% fewer parameters) maintains accuracy. The iTransformer baseline retains a statistically significant advantage over PowerMamba (), reflecting the trade-off between Mamba’s linear scalability and Transformer’s pairwise correlation modeling.
@@ 6 Results
6.1 Multi-Horizon Performance (walk-forward)
Evaluation regime. This subsection reports results from the rolling-origin walk-forward backtest described in Section 5.3.
Table 1 presents multi-horizon MAPE. PowerMamba is the strongest Mamba variant at the 1-hour horizon, while iTransformer achieves the lowest MAPE at longer horizons (6h–24h) in our walk-forward evaluation. These accuracy rankings, however, do not fully determine operational tail risk (Reserve), motivating the grid-specific metrics reported later.
@@ 6.1 Multi-Horizon Performance (walk-forward)
| S-Mamba | 16.4M | 5.49 0.12 | 3.86 0.02 | 3.40 0.47 | 8.03 0.18 |
| Mamba-ProbTSF | 16.4M | 5.23 0.04 | 3.91 0.19 | 3.30 0.08 | 7.89 0.17 |
| PowerMamba | 2.5M | 4.44 0.20 | 4.10 0.09 | 3.62 0.30 | 7.31 0.06 |
| LSTM | 0.21M | 7.85 0.20 | 5.24 0.13 | 9.46 0.32 | 9.99 0.13 |
| iTransformer | 50.0M | 4.64 0.10 | 3.36 0.05 | 3.10 0.21 | 6.85 0.08 |
| Chronos (zero-shot)† | 8.0M | 2.10 | 3.21 | 3.72 | 7.59 |
| †Zero-shot evaluation; no training or uncertainty estimation. | |||||
6.2 Error Analysis: The Impact of Weather on Forecast Reliability (walk-forward)
Evaluation regime. This subsection uses the walk-forward backtest forecasts (all lead times within the 48h horizon) described in Section 5.3.
The integration of weather data into deep learning models for load forecasting has been extensively studied [1, 5]. Prior work has demonstrated improvements using LSTM-based architectures with temperature features [22], CNN-GRU hybrids for extreme weather scenarios, and attention-based models for cross-variable correlation. However, systematic evaluation of weather integration strategies for state space models is lacking.
@@ 6.2 Error Analysis: The Impact of Weather on Forecast Reliability (walk-forward)
6.3 Weather Integration Results (walk-forward)
Evaluation regime. This subsection reports walk-forward results on matched evaluation windows (Section 5.3).
@@ 6.3 Weather Integration Results (walk-forward)
Overall, weather integration materially narrows the error distribution in the regimes where temperature-driven load spikes dominate. Next, we test whether probabilistic calibration and explicit Bias/OPR constraints can further reduce tail events without allowing trivial “safety” via inflation.
6.4 Probabilistic Calibration and Bias/OPR-Constrained Objectives (fixed split)
Evaluation regime. This subsection reports fixed-split fine-tuning results for loss-function ablations (Section 5.6); it is not directly comparable to the walk-forward results above.
Loss-function hypothesis. While weather integration improves robustness, error distributions still exhibit rare large deviations. We hypothesized that some tail events are loss-driven: standard MSE training treats all errors symmetrically and does not target calibrated upper-tail behavior.
Multi-quantile pinball experiment. We fine-tune with a weighted multi-quantile pinball objective (Eq. 15). On the CA ISO-TAC fixed evaluation split (24h), this reduces large-error events (e.g., 1000 MW: 1,111 917; 2000 MW: 365 241) while also improving MAPE (Fig. 4). However, these gains can reflect forecast inflation; we therefore report and (when needed) constrain Bias24h and Over-Prediction Rate (OPR) alongside tail metrics.
Table 2 presents an operator-legible breakdown (CA ISO-TAC, 24h), explicitly reporting Bias24h and OPR alongside tail-risk Reserve. Three high-signal observations:
Tail metrics can improve via inflation. For iTransformer, WeatherMulti-Q reduces Reserve (28.70% 13.83%) but does so with large positive bias (+1,862 MW) and high OPR (78.8%), indicating a shift in scheduled location.
Constraints enforce auditable trade-offs. Adding Bias/OPR control reduces inflation for iTransformer (Bias24h +1,862 +456 MW; OPR 78.8% 61.6%) while partially trading off tail-risk (Reserve 13.83% 15.18%).
Constraint tuning is architecture-dependent. Some architectures exhibit unstable behavior under the same constraint settings (e.g., PowerMamba inflates under BiasCtrl), reinforcing that risk-aversion should be posed as constrained optimization with auditable budgets rather than a single universal loss.
| Model | Loss | MAPE | UPR | Reserve | Bias24h (MW) | OPR |
| S-Mamba | MSE | 4.42% | 43.5% | 16.28% | +243.6 | 56.5% |
| Multi-Q | 3.63% | 34.0% | 13.65% | +405.4 | 66.0% | |
| +BiasCtrl | 3.86% | 31.2% | 12.75% | +507.6 | 68.8% | |
| PowerMamba | MSE | 3.95% | 38.4% | 12.91% | +370.9 | 61.6% |
| Multi-Q | 4.19% | 34.6% | 13.65% | +480.1 | 65.4% | |
| +BiasCtrl | 9.01% | 6.2% | 7.42% | +2027.1 | 93.8% | |
| Mamba-ProbTSF | MSE | 4.26% | 44.9% | 14.86% | +176.5 | 55.1% |
| Multi-Q | 3.69% | 36.2% | 14.17% | +355.2 | 63.8% | |
| +BiasCtrl | 3.84% | 64.3% | 17.45% | -325.9 | 35.7% | |
| iTransformer | MSE | 9.34% | 50.6% | 28.70% | +109.3 | 49.4% |
| Multi-Q | 10.08% | 21.2% | 13.83% | +1862.3 | 78.8% | |
| +BiasCtrl | 5.04% | 38.4% | 15.18% | +456.4 | 61.6% |
| Model | Variant | Bias24h (MW) | Reserve | UPR (%) | OPR (%) |
| S-Mamba | MSE | +243.6 | 16.28 | 43.5 | 56.5 |
| Multi-Q | +405.4 | 13.65 | 34.0 | 66.0 | |
| +BiasCtrl | +507.6 | 12.75 | 31.2 | 68.8 | |
| PowerMamba | MSE | +370.9 | 12.91 | 38.4 | 61.6 |
| Multi-Q | +480.1 | 13.65 | 34.6 | 65.4 | |
| +BiasCtrl | +2027.1 | 7.42 | 6.2 | 93.8 | |
| Mamba- ProbTSF | MSE | +176.5 | 14.86 | 44.9 | 55.1 |
| Multi-Q | +355.2 | 14.17 | 36.2 | 63.8 | |
| +BiasCtrl | -325.9 | 17.45 | 64.3 | 35.7 | |
| iTransformer | MSE | +109.3 | 28.70 | 50.6 | 49.4 |
| Multi-Q | +1862.3 | 13.83 | 21.2 | 78.8 | |
| +BiasCtrl | +456.4 | 15.18 | 38.4 | 61.6 |
Effect of probabilistic calibration on tail events. Figure 5 compares (i) Weather MSE, (ii) WeatherMulti-Q, and (iii) WeatherMulti-Q+BiasCtrl on a single consistent evaluation split (same targets across variants). Multi-quantile calibration reduces large-error counts relative to Weather MSE, while adding Bias/OPR constraints partially trades off tail improvements for reduced inflation risk (Table 2).
@@ 6.5 NEMs (BTM Registry) Integration Results
Evaluation regime. This subsection reports walk-forward results on utilities where NEMs/registry features are available. We next evaluate whether static NEMs/registry-derived BTM features (e.g., installed PV and storage capacity) provide incremental predictive value on top of the weather-aligned baseline. This analysis is reported only for utilities where NEMs/registry features were available in our pipeline (PGE, SCE, SDGE, TIDC).
@@
Figure 7 summarizes the incremental impact of NEMs integration relative to the weather baseline, measured as changes in MAPE and Reserve (tail under-forecast requirement proxy). Improvements are heterogeneous: in the PV-dominant reference utility with NEMs availability (PGE), NEMs integration reduces both average error and tail-risk reserve requirement. In contrast, other territories show smaller changes and one (SCE) degrades slightly. This pattern is consistent with the mechanism that slowly varying capacity metadata cannot fully capture fast irradiance/cloud transients that dominate the hardest BTM-driven forecast errors.
@@ 7.1 Performance Comparison with CAISO Benchmarks
| iTransformer | 5.37% | +18.0% | 50.0M | No |
| Chronos (zero-shot) | 7.59% | +66.8% | 8.0M | No |
@@
7.2 Temperature, Tail Risk, and Operator Implications
Our error analysis shows a statistically significant but modest association between temperature and forecast error magnitude (), supporting the operational intuition that weather is a relevant covariate and that loss-function modifications cannot substitute for missing exogenous information. More importantly for operations, the walk-forward results reinforce that reserve requirements are not determined by MAPE alone: models with similar average accuracy can imply very different one-sided tail requirements (Reserve) under under-forecast risk. Weather integration narrows the error distribution across utilities (Fig. 3), which can translate directly into reduced upward reserve needs during temperature-driven demand spikes.
@@ 7.4 Foundation Model Performance
Chronos-T5-Base (200M parameters) performed no better than Chronos-Small (8M) in zero-shot evaluation (e.g., 5.76% vs 5.23% MAPE at the 24-hour horizon). This suggests that foundation model scaling does not automatically translate to domain-specific performance. General-purpose pre-training lacks the structural priors needed to handle the specific physical constraints of the grid—such as the complex interaction between thermal inertia and behind-the-meter generation—reinforcing the value of specialized architectures and weather integration strategies.
7.5 Limitations from Unobserved BTM Generation
A recurring theme in our results is the ceiling imposed by data observability. Our "grey box" experiments with static NEMs registry features (Section 6.5) yielded only marginal gains. Figure 7 shows that even with capacity data, the model struggles to capture the dynamic, weather-dependent ramping of distributed solar. This points to a fundamental limitation: net load signals conflate grid demand with behind-the-meter (BTM) generation, and static capacity metadata is insufficient to disentangle them.
This "visibility gap" suggests that future improvements will not come from larger black-box transformers, but from structural end-to-end learning. As demonstrated by Shi et al. [17], differentiable optimization layers can explicitly model physical relationships (such as solar generation physics or price response) within the neural network. By adopting such end-to-end decomposition methods to dynamically learn latent BTM states from aggregate net load, future work could address the visibility gap that static feature engineering failed to close.
@@ 7.5 Limitations from Unobserved BTM Generation
Baseline Coverage. Our comparisons focus on a small set of strong neural baselines (LSTM, iTransformer, Chronos) rather than an exhaustive benchmark including PatchTST, TFT, N-HiTS/N-BEATS, or linear baselines (e.g., DLinear/NLinear). Future work should include autocorrelation-aware significance testing for a more complete statistical analysis.
@@ 8 Conclusion
Mamba-based state space models achieve competitive accuracy for California grid load forecasting. With weather integration, PowerMamba achieves 4.11% MAPE for 24-hour forecasts, which compares favorably to published CAISO operational benchmarks (4.55% MAPE; contextual reference). Critically, our walk-forward evaluation demonstrates that operational tail risk (Reserve) can diverge materially from average accuracy, motivating grid-specific reliability metrics alongside MAPE for safety-critical deployment decisions.
@@
Probabilistic calibration must be paired with bias controls. Multi-quantile objectives can reduce large-error events and tail reserve requirements, but can also be “gamed” by systematic inflation if Bias/OPR are not monitored and constrained. Our Bias/OPR-constrained objective provides an operator-legible way to enforce auditable trade-offs between tail-risk and schedule inflation (Table 2, Table 3). Operational Deployment Implications. Reduced tail under-forecast risk (Reserve) can translate to lower upward reserve requirements and reduced reliance on fast-ramping resources. The computational efficiency of Mamba architectures enables real-time deployment without significant infrastructure upgrades, while calibrated uncertainty quantification from Mamba-ProbTSF supports probabilistic reserve scheduling.
@@ Appendix C Multi-Seed Robustness
To assess robustness to random initialization, we trained PowerMamba + Weather with seeds . The 24-hour MAPE on CA ISO-TAC was (mean std), confirming that reported results are stable across initializations. Full multi-seed results for all architectures will be provided in the final version.
@@ Appendix D Per-Utility Results
Smaller utilities (SDGE-TAC, TIDC) exhibit higher MAPE due to increased load volatility relative to system size. Weather integration tends to improve accuracy, with the largest relative gains in smaller, more volatile territories; however, improvements are not uniform across all regions (Table 7).